OpenStudy (anonymous):

express as a single fraction the exact value of sin 75

OpenStudy (mertsj):

\[\sin75=\sin(30+45)=\sin30\cos45+\cos30\sin45\]

OpenStudy (anonymous):

sin75=sin(45+30)=sin45.cos30+cos45.sin30=(\[\sqrt{3}/2+1/2\])\[\div\]\[\sqrt{2}\] =(sqrt(3)+1)/(2*sqrt(2)).

OpenStudy (mertsj):

\[=\frac{1}{2}(\frac{\sqrt{2}}{2})+\frac{\sqrt{3}}{2}(\frac{\sqrt{2}}{2})=\frac{\sqrt{2}+\sqrt{6}}{4}\]